OpenStudy (anonymous):

Juliana opened a new escort business. She hired three girls: Destiny, Majestic and Heaven. All of Juliana's employees are allowed to charge their clients any rate of their choosing, but they must give Juliana 50% of their income. This is to pay for drivers and protection. Destiny charges $300 per hour, Majestic is pretty nasty, so she only charges $30 per hour, and Heaven charges $500 per hour. If all three girls worked one hour on their first day, how much did Juliana make from the girls?

OpenStudy (anonymous):

This is homework btw. I don't get it.

OpenStudy (bibby):

Ignoring those exorbitant rates, You add up the amount of money they all made in an hour of working and multiply by 50% or 0.50

OpenStudy (bibby):

seriously who pays 300 bucks for an escort

OpenStudy (anonymous):

.5(300) + .5(30) + .5(500) = ?

OpenStudy (anonymous):

wat is a escort

OpenStudy (anonymous):

400?

OpenStudy (anonymous):

what

OpenStudy (bibby):

Prostitutes provide sex/sexual favors for money. Escorts provide companionship and romance (and ???) for money.

OpenStudy (the_fizicx99):

^

OpenStudy (anonymous):

omg

OpenStudy (anonymous):

who wrote this book lol

OpenStudy (anonymous):

the 2nd part of the problem is write a linear equation with the variable x representing juliana's cut bc juliana might change it day to day

OpenStudy (anonymous):

@tHe_FiZiCx99 @bibby @mathbrz

OpenStudy (anonymous):

0.5x + 0.5x + 0.5x ?

OpenStudy (anonymous):

but what is it equal too

OpenStudy (anonymous):

300x + 30x + 500x = j

OpenStudy (anonymous):

thx

OpenStudy (doc.brown):

It distributes too, if you're interested. I mean the numbers, not STDs. \[300x+30x+500x=j\]\[x(300+30+500)=j\]\[j=1100x\]

OpenStudy (elisaneedshelp):

this is a crazy nasty problem